# Topological Phases of Many-Body Localized Systems: Beyond Eigenstate Order

### v2

@@ A.3 Non-uniqueness
A QCA which maps $H$ to a sum of product of $\sigma^{z}_{k}$ operators always exists, but this QCA is not unique.
Indeed, the QCA $V$ is a unitary operator which diagonalizes $H$ in the simultaneous eigenbasis of $\{\sigma^{z}_{k}\}_{k\in K}$. Even if $H$ is nondegenerate, such unitaries are only unique up to reordering of eigenvalues and multiplying eigenstates by arbitrary phases. Specifically, if $P$ is a unitary generalized permutation matrix (a unitary matrix with exactly one unit modulus complex number in each row and column) in the eigenbasis of $\{\sigma^{z}_{k}\}_{k\in K}$, then $(VP)^{\dagger}H(VP)$ is still diagonal. If $H$ is nondegenerate, the full redundancy of the correspondence between localized models with short-ranged entanglement and QCA is found by restricting $P$ to be a QCA.
@@
Mathematically, the classification of localized phases with short-ranged entanglement is given by a quotient of $\pi_{0}(\mathrm{QCA}_{d})$. Denoting the set of generalized permutation QCA by $\mathrm{GP}_{d}$, there is a canonical inclusion map $\iota:\mathrm{GP}_{d}\to\mathrm{QCA}_{d}$, and an induced homomorphism $\iota^{*}:\pi_{0}(\mathrm{GP}_{d})\to\pi_{0}(\mathrm{QCA}_{d})$. We write the image of this homomorphism as $I_{0}=\iota^{*}(\pi_{0}(\mathrm{GP}_{d}))$. Then the quotient which classifies localized topological phases is
$\pi_{0}(\mathrm{QCA}_{d})/I_{0}.$ (37)
@@
[…]
@@
and $\pi_{0}(\mathrm{QCA}_{d}/\mathrm{GP}_{d})\cong\pi_{0}(\mathrm{QCA}_{d})/I_{0}$ follows from the first isomorphism theorem applied to $q^{*}$.
@@
The only example for which this redundancy affects the classification of localized phases for $d\leq 3$ is the classification of Floquet ($n=1$) phases in zero dimensions [31]. Looking ahead to Appendix C.2, we have that $\pi_{1}(\mathrm{QCA}_{0})={\mathbb{Q}}/{\mathbb{Z}}$, and that a complete index is given by the winding number $\tilde{w}_{1}[V]$. In zero dimensions, all unitaries are QCA, so the diagonal unitary with elements
@@
$\tilde{w}_{1}[VP_{m}]=\tilde{w}_{1}[V]+\frac{m}{N}\mod 1,$ (45)
@@
where $N$ is the Hilbert space dimension. Thus, multiplication by $P_{m}(\theta)$ can take $\tilde{w}_{1}[V^{\prime}]$ to any allowed value, and we conclude that all loops of zero dimensional QCA (loops of projective unitaries) define trivial localized phases.
@@
More systematically, we can again classify driven localized phases by a quotient of the group $\pi_{0}\left[{\mathbb{T}}^{n},\mathrm{QCA}_{d}\right]$. The generalized permutations $P({\vec{\theta}})$ are particular maps from ${\mathbb{T}}^{n}$ to $\mathrm{QCA}_{d}$, and we denote the subgroup of $\pi_{0}\left[{\mathbb{T}}^{n},\mathrm{QCA}_{d}\right]$ induced by the inclusion of the generalized permutations in the full space of maps from ${\mathbb{T}}^{n}$ to $\mathrm{QCA}_{d}$ by $I_{n}$. Then the classification of $n$-tone driven localized phases in $d$-dimensions with short-ranged entanglement is given by
$\pi_{0}\left[{\mathbb{T}}^{n},\mathrm{QCA}_{d}\right]/I_{n}.$ (46)
@@
This group is not isomorphic to $\pi_{0}\left[{\mathbb{T}}^{n},\mathrm{QCA}_{d}/\mathrm{GP}_{d}\right]$ in general. Taking, for instance, $n=1$, there are loops in $\pi_{0}\left[{\mathbb{T}}^{1},\mathrm{QCA}_{d}/\mathrm{GP}_{d}\right]\cong\pi_{1}(\mathrm{QCA}_{d}/\mathrm{GP}_{d})$ which correspond to paths from the identity to a locally generated permutation. This is not a legitimate micromotion operator $V(\theta)$, as it is not periodic as a (projective) unitary.
@@ Appendix B The $\Omega$-spectrum property
In topology, an $\Omega$-spectrum refers to a sequence of topological spaces $(Y_{d})_{d\in{\mathbb{Z}}}$ together with (weak) homotopy equivalences
$Y_{d-1}\simeq\Omega Y_{d}$ (47)
where $\Omega Y_{d}$ is the space of loops in $Y_{d}$ which begin and end at a specified base point. $\Omega$-spectra play a special role in topology, as they serve to define generalized cohomology theories—functors from the category of topological spaces to abelian groups obeying certain axioms [95, 102, 104, 103]. The functor in question is
$h^{d}_{Y}(X)=\pi_{0}\left[X,Y_{d}\right].$ (48)
The $\Omega$-spectrum property comes with substantial structure, and highly constrains the properties of the spaces $Y_{d}$.
@@
For $d<0$, we define $\mathrm{QCA}_{d}=\Omega^{-d}\mathrm{QCA}_{0}$, making the negative degrees of this statement trivial. In positive degree, we propose that the swindle map $S$ [Eq. (8)] provides the homotopy equivalence. Thus, we propose there is a generalized cohomology theory
$h^{d}(X)=\pi_{0}\left[X,\mathrm{QCA}_{d}\right].$ (50)
In this appendix, we provide arguments in support of this conjecture. In Appendix B.1, we recall that if loops of QCA can be put on open boundary conditions such that they act by QCA on their edges—known as blending with the identity—then loops of QCA can all be parameterized as finite depth circuits. We also argue that this condition holds for loops of one-dimensional QCA. Then, in Appendix B.2, we construct an explicit deformation retract from the space of circuits to the image of the swindle map. As the swindle map is injective, this shows
@@
under the assumptions of Appendix B.1.
Making these arguments rigorous requires being more specific about the topology on $\mathrm{QCA}_{d}$, even in $d=1$. This can be done, but we defer such technicalities to Ref. [107].
@@ B.1 Blending and circuits
A blending between $\alpha$ and $\beta\in\mathrm{QCA}_{d}$ is a third QCA $\gamma$ such that
@@
where $\mathrm{supp}(A)$ is the smallest subset of ${\mathbb{Z}}^{d}$ on which the operator $A$ acts non-trivially, and $x_{l,r}$ are positions along the first axis in ${\mathbb{Z}}^{d}$. That is, $\gamma$ agrees with $\alpha$ to the left of some cut, and agrees with $\beta$ to the right of the cut. We can similarly define a blending $\gamma({\vec{\theta}})$ between parameterized families of QCA $\alpha({\vec{\theta}})$ and $\beta({\vec{\theta}})$. Two (families of) QCA are said to blend into each other if a blending exists between them.
Two QCA blend into one another if and only if they are related by a finite depth circuit [94]. That is, $\alpha$ blends with $\beta$ at any position of the cut $x$ if and only if $\alpha\beta^{-1}$ is equal to a finite depth unitary circuit. (Actually, it is sufficient for a blending to exist at any position [109].)
@@
We can split $\gamma$ into a depth two circuit by iteratively splitting off gates of the form $\tilde{\gamma}_{x}\tilde{\gamma}_{x+1}^{-1}$ or $\tilde{\gamma}^{-1}_{x}\tilde{\gamma}_{x+1}$. (The gates are themselves QCA in $d-1$ dimensions.) First, observe that we can split $\gamma$ into a product of two semi-infinite QCA
@@
In Ref. [107] we study this structure in greater detail, and characterize the circuit liftings implicit here as a fibration on the space of QCA.
@@ B.2 Deformation to the swindle
where the second equality comes from translating Eq. (59) by one unit towards positive $x$ and inverting the equation. We see that $u_{0,1}(v,t)^{2}=u_{0,1}(w,t)^{2}$. Making a spectral decomposition of both sides, we conclude that $u_{0,1}(v,t)=u_{0,1}(w,t)$ up to multiplying eigenvalues by ±$\pm 1$. However, these paths both start at $t=0$ with the same eigenvalues $\lambda=+1$, which remain nonzero throughout all $t$. Thus, all the discrete choices of ±$\pm 1$ must be $+1$, and $u_{0,1}(v,t)=u_{0,1}(w,t)$. Then comparing these at $t=1$ gives $v_{0}v^{\dagger}_{1}=w_{0}w^{\dagger}_{1}$, and hence $v=w$ up to a phase. We regard QCA as projective unitaries, so $v=w$.
@@
Figure 4: The loop of finite depth circuits $\gamma(t)\otimes S(v_{0}^{\gamma},s)^{\dagger}$ is deformed to the constant loop at the identity as a function of $s\in[0,1]$. (a) At $s=0$, the circuit consists of two layers. Between $s=0$ and $s=1/2$ the gates $u^{\dagger}_{x,x+1}(v_{0}^{\gamma},t)$ in the top layer (red) are swapped to the bottom layer while having their order of multiplication reversed. (b) The gate $u^{\dagger}_{0,1}(v_{0}^{\gamma},t)w_{0,1}(t)$ [$u^{\dagger}_{-1,0}(v_{0}^{\gamma},t)w_{-1,0}(t)$] acts as the identity on site 00 for $t\in\{0,1\}$. It can thus be swapped (over $s\in[\tfrac{1}{2},\tfrac{3}{4}]$) so that it acts on site $1$ [$-1$] and its ancilla while ensuring that the whole circuit is still a loop. (c) Between $s=\tfrac{3}{4}$ and $s=1$ the circuit (red) is “rolled up”, such that it acts as the identity (gray dots) in any finite region.
@@
$w^{\prime}_{0,1}(1)=\mathbbm{1}\otimes v^{\gamma}_{0}v^{\gamma\dagger}_{1}\text{ and }w^{\prime}_{-1,0}(1)=v^{\gamma\dagger}_{0}v^{\gamma}_{-1}\otimes\mathbbm{1}$ (67)
both act as the identity on site $x=0$ at $t=1$. Thus, in $s\in[\tfrac{1}{2},\tfrac{3}{4}]$, we can introduce $\mathrm{SWAP}(4(s-\tfrac{1}{2}))$ gates to the positions shown in Fig. 4(b) which “roll up” the circuit—the gates previously acting on site 00 now act on-site at $x=\pm 1$ between ancillae. Explicitly, for $s\in[\tfrac{1}{2},\tfrac{3}{4})$, the gate in $E(\gamma,t,s)$ acting across the $(00^{\prime},11^{\prime})$ edge is
@@
As $w^{\prime}_{0,1}(1)$ acts as the identity on site 0, the introduced swap gates cancel at $t=1$, and the entire circuit remains a loop for all $s$. By $s=\tfrac{3}{4}$, the gate acts as the identity on site 0 for all $t\in[0,1]$. We repeat an analogous deformation on the other gate with support on site 00. The part of the circuit which is supported on $x=0$ is thus rolled up to act on the ancillae at sites $x=\pm 1$.
This rolling up procedure can be iterated. At $s=\frac{3}{4}$, we see that the gate $w^{\prime}_{1,2}(t)w^{\prime}_{1^{\prime},1}(t)$ is the identity on sites $1$ and $1^{\prime}$ when $t=1$. Thus, by adding ancillae to site $2$, this gate can also be rolled up within $s\in[\tfrac{3}{4},\tfrac{7}{8}]$ [Fig. 4(c)]. We continue rolling up sites $x$ and $-x$ within $s\in[1-2^{-x-1},1-2^{-x-2}]$, such that, as $s\to 1$, we have $E(\gamma,t,s)\to\mathbbm{1}\otimes S(v_{0}^{\gamma},t)$ on any finite patch of the lattice. Technically, this makes $E(\gamma,t,s)$ continuous in the strong operator topology.

### v3

@@ A.3 Non-uniqueness
A QCA which maps $H$ to a sum of product of $\sigma^{z}_{k}$ operators always exists, but this QCA is not unique. Here, we characterize this redundancy, and show that it does not affect the classification of driven short-range entangled MBL phases unless $(d,n)=(0,1)$.
Indeed, diagonalization of the Hamiltonian $H=VH_{\mathrm{triv}}V^{\dagger}$ (where $H_{\mathrm{triv}}$ is a sum of products of $\sigma^{z}_{k}$ operators) only determines $V$ up to right-multiplication by a generalized permutation QCA. We define a QCA $P$ to be a generalized permutation exactly when it maps $\sigma^{z}_{k}$ operators to sums of products of other $\sigma^{z}_{k}$s. This name is motivated as follows. In a finite dimensional Hilbert space, $V$ is a unitary operator which diagonalizes $H$ in the simultaneous eigenbasis of $\{\sigma^{z}_{k}\}_{k\in K}$. Even if $H$ is nondegenerate, such unitaries are only unique up to reordering of eigenvalues and multiplying eigenstates by arbitrary phases. Specifically, if $P$ is a unitary generalized permutation matrix (a unitary matrix with exactly one unit modulus complex number in each row and column) in the eigenbasis of $\{\sigma^{z}_{k}\}_{k\in K}$, then $(VP)^{\dagger}H(VP)$ is still diagonal. We extend this “generalized permutation” language to the infinite dimensional case via the property of preserving the algebra of $\sigma^{z}_{k}$ operators.
@@
Mathematically, the classification of localized phases with short-ranged entanglement is given by a quotient of $\pi_{0}(\mathrm{QCA}_{d})$. Denoting the set of generalized permutation QCA by $\mathrm{GP}_{d}$, there is a canonical inclusion map $\iota:\mathrm{GP}_{d}\to\mathrm{QCA}_{d}$, and an induced homomorphism $\iota^{*}:\pi_{0}(\mathrm{GP}_{d})\to\pi_{0}(\mathrm{QCA}_{d})$. We write the image of this homomorphism as $I_{d,0}=\iota^{*}(\pi_{0}(\mathrm{GP}_{d}))$. Then the quotient which classifies localized topological phases is
$\pi_{0}(\mathrm{QCA}_{d})/I_{d,0}.$ (135)
@@
$\mathrm{GP}_{d}\xrightarrow{\iota}\mathrm{QCA}_{d}\xrightarrow{q}\mathrm{QCA}_{d}/\mathrm{GP}_{d}$ (136)
@@
and $\pi_{0}(\mathrm{QCA}_{d}/\mathrm{GP}_{d})\cong\pi_{0}(\mathrm{QCA}_{d})/I_{d,0}$ follows from the first isomorphism theorem applied to $q^{*}$.
@@
The only example for which this redundancy affects the classification of driven localized phases for $d\leq 3$ is the classification of Floquet ($n=1$) phases in zero dimensions. From Sec. III, we have that $\pi_{1}(\mathrm{QCA}_{0})\cong{\mathbb{Q}}/{\mathbb{Z}}$, and that a complete index is given by the winding number $\widetilde{W}_{1}[V]$. In zero dimensions, all unitaries are QCA, so the diagonal unitary with elements
@@
$\widetilde{W}_{1}[VP_{m}]=\widetilde{W}_{1}[V]+\frac{m}{N}\mod 1,$ (143)
@@
where $N$ is the Hilbert space dimension. Thus, multiplication by $P_{m}(\theta)$ can take $\widetilde{W}_{1}[V^{\prime}]$ to any allowed value, and we conclude that all loops of zero dimensional QCA (loops of projective unitaries) define trivial localized phases.
@@
To see that the redundancy by generalized permutations does not affect the driven classification for other value of $(d,n)$, we must be more systematic. We can classify driven localized phases by a quotient of the group $\pi_{0}\left[{\mathbb{T}}^{n},\mathrm{QCA}_{d}\right]$, where ${\mathbb{T}}^{n}=[{\mathbb{R}}/(2\pi{\mathbb{Z}})]^{n}$ is the $n$-torus. The generalized permutations $P({\vec{\theta}})$ are particular maps from ${\mathbb{T}}^{n}$ to $\mathrm{QCA}_{d}$, and we denote the subgroup of $\pi_{0}\left[{\mathbb{T}}^{n},\mathrm{QCA}_{d}\right]$ that includes such maps by $I_{d,n}$. Then the classification of $n$-tone driven localized phases in $d$-dimensions with short-ranged entanglement is given by
$\pi_{0}\left[{\mathbb{T}}^{n},\mathrm{QCA}_{d}\right]/I_{d,n}.$ (144)
@@
Note that this group is not isomorphic to $\pi_{0}\left[{\mathbb{T}}^{n},\mathrm{QCA}_{d}/\mathrm{GP}_{d}\right]$ in general. Taking, for instance, $n=1$, there are loops in $\pi_{0}\left[{\mathbb{T}}^{1},\mathrm{QCA}_{d}/\mathrm{GP}_{d}\right]\cong\pi_{1}(\mathrm{QCA}_{d}/\mathrm{GP}_{d})$ which correspond to paths from the identity to a locally generated permutation. This is not a legitimate micromotion operator $V(\theta)$, as it is not periodic as a (projective) unitary.
Assuming that a generalized permutation $P({\vec{\theta}})$ acts as a family of finite depth circuits, we show that the homotopy class of $P({\vec{\theta}})$ is trivial unless $(d,n)=(0,1)$. This implies that $I_{d,n}\leq\pi_{0}(\mathrm{QCA}_{d})$ for $(d,n)\neq(0,1)$, and we suspect that $I_{d,n}\cong 0$ unless $d+n=1$, so that the only redundancy in the description of MBL phases by QCA phases is for $(d,n)\in\{(0,1),(1,0)\}$.
As we assume $P({\vec{\theta}})$ acts by finite depth circuits, we only need to consider $n\geq 1$ In fact, Ref. [31] already showed that $I_{0,n}\cong 0$ for $n\neq 1$, so we can also restrict to $d\geq 1$. In this case, we argue that $P({\vec{\theta}})$ can always be constructed as a depth-two circuit of (d1)(d-1)-dimensional generalized permutations. Such a circuit does not pump any nontrivial QCA to its boundary on open boundaries, so if the bulk-boundary correspondence discussed in Sec. III holds, then all such circuits are trivial. More constructively, the deformation discussed in Appendix B.2 can be applied to such a circuit to explicitly give a homotopy to the identity.
To arrive at this depth-two circuit, we observe that, by definition, any generalized permutation $P({\vec{\theta}})$ is generated by time evolution under a time-independent Hamiltonian composed of $\sigma^{z}$ LIOMs. That is, there are trivial LIOM Hamiltonians $H_{j}$ such that
$i\partial_{\theta_{j}}P({\vec{\theta}})]P({\vec{\theta}})^{\dagger}=H_{j}.$ (145)
All the $H_{j}$ operators commute, so we can solve this equation as $P({\vec{\theta}})=e^{-i{\vec{\theta}}\cdot\vec{H}}P(0)$, where $\vec{H}=(H_{1},...,H_{n})$. Split $P(0)$ into a brickwork circuit of (d1)(d-1)-dimensional blocks
$P(0)=\left(\prod_{i\in 2{\mathbb{Z}}+1}b_{i}^{\dagger}\right)\left(\prod_{i\in 2{\mathbb{Z}}}b_{i}\right),$ (146)
where each $b_{i}$ only overlaps in support with $b_{i\pm 1}$ near their boundaries, and such that the width of these blocks are much larger than the support of any term in $H_{j}$. We want to restrict the Hamiltonians $H_{j}$ to $H^{(2i)}_{j}$ with support contained within that of $b_{2i}$, such that $H^{(2i)}_{j}$ agrees with $H_{j}$ near the center of the block, and such that $e^{-i{\vec{\theta}}\cdot\vec{H}^{(2i)}}$ is still periodic. To construct $H^{(2i)}_{j}$, first define $\tilde{H}^{(2i)}_{j}$ to include every term in $H_{j}$ with support overlapping the $b_{2i}$ block. Then pick an eigenvalue $\lambda_{k}$ of each $\sigma^{z}_{k}$ with $k$ not inside the $b_{2i}$ block, and make a replacement
$\sigma^{z}_{k_{1}}\cdots\sigma^{z}_{k_{l}}\otimes A\mapsto\lambda_{k_{1}}\cdots\lambda_{k_{l}}\mathbbm{1}\otimes A$ (147)
[…]
$P({\vec{\theta}})\left(\prod_{i\in 2{\mathbb{Z}}}b_{i}^{\dagger}e^{i{\vec{\theta}}\cdot\vec{H}^{(i)}}\right)=\prod_{i\in 2{\mathbb{Z}}+1}b_{i}^{\dagger}({\vec{\theta}})$ (148)
must split into a tensor product of disjoint odd-indexed blocks, as $P({\vec{\theta}})$ and $e^{-i{\vec{\theta}}\cdot\vec{H}^{(2i)}}b_{2i}$ agree deep inside the even-indexed blocks. As products of generalized permutaions, these blocks are also generalized permutations. Thus, we have the desired brickwork circuit for $P({\vec{\theta}})$, and conclude that $P({\vec{\theta}})$ is homotopic to the identity as a map from ${\mathbb{T}}^{n\geq 1}$ to $\mathrm{QCA}_{d\geq 1}$.
@@ Appendix B The $\Omega$-spectrum property
For $d<0$, we define $\mathrm{QCA}_{d}=\Omega^{-d}\mathrm{QCA}_{0}$, making the negative degrees of this statement trivial. In positive degree, we propose that the swindle map $S$ (Fig. 3) provides the homotopy equivalence.
In this appendix, we provide further arguments in support of this conjecture. In Appendix B.1, we recall that if loops of QCA can be put on open boundary conditions such that they act by QCA on their edges—known as blending with the identity—then loops of QCA can all be parameterized as finite depth circuits of lower-dimensional QCA. We also argue that this condition holds for loops of one-dimensional QCA. Then, in Appendix B.2, we construct an explicit deformation retract from the space of circuits to the image of the swindle map. As the swindle map is injective, this shows
@@
under the assumptions of Appendix B.1, where $\cong$” is homeomorphism.
[…]
@@ B.1 Blending and circuits
A blend between $\alpha$ and $\beta\in\mathrm{QCA}_{d}$ is a third QCA $\gamma$ such that
@@
where $\mathrm{supp}(A)$ is the smallest subset of ${\mathbb{Z}}^{d}$ on which the operator $A$ acts non-trivially, $x_{l,r}$ are positions along the first axis in ${\mathbb{Z}}^{d}$, and $\alpha(\cdot)$ is the map on local operators defined by conjugation by the locality-preserving unitary defining $\alpha$. That is, $\gamma$ agrees with $\alpha$ to the left of some cut, and agrees with $\beta$ to the right of the cut. We can similarly define a blend $\gamma({\vec{\theta}})$ between parameterized families of QCA $\alpha({\vec{\theta}})$ and $\beta({\vec{\theta}})$. Two (families of) QCA are said to blend into each other if a blend exists between them.
Two QCA blend into one another if and only if they are related by a finite depth circuit of lower-dimensional QCAs [99]. That is, $\alpha$ blends with $\beta$ at any position of the cut $x$ if and only if $\alpha\beta^{-1}$ is equal to a finite depth unitary circuit. (Actually, it is sufficient for a blend to exist at any position [110].)
@@
We can split $\gamma$ into a depth two circuit by iteratively splitting off gates of the form $\tilde{\gamma}_{x}\tilde{\gamma}_{x+1}^{-1}$ or $\tilde{\gamma}^{-1}_{x}\tilde{\gamma}_{x+1}$. (The gates are themselves QCA in d1d-1 dimensions.) First, observe that we can split $\gamma$ into a product of two semi-infinite QCA
@@ B.2 Deformation to the swindle
where the second equality comes from translating Eq. (158) by one unit towards positive $x$ and inverting the equation. We see that $u_{0,1}(v,t)^{2}=u_{0,1}(w,t)^{2}$. Making a spectral decomposition of both sides, we conclude that $u_{0,1}(v,t)=u_{0,1}(w,t)$ up to multiplying eigenvalues by $\pm 1$. However, these paths both start at $t=0$ with the same eigenvalues $\lambda=+1$, which remain nonzero throughout all $t$. Thus, all the discrete choices of $\pm 1$ must be +1+1, and $u_{0,1}(v,t)=u_{0,1}(w,t)$. Then comparing these at $t=1$ gives $v_{0}v^{\dagger}_{1}=w_{0}w^{\dagger}_{1}$, and hence $v=w$ up to a phase. We regard QCA as projective unitaries, so $v=w$.
@@
Figure 6: The loop of finite depth circuits $\gamma(t)\otimes S(v_{0}^{\gamma},s)^{\dagger}$ is deformed to the constant loop at the identity as a function of $s\in[0,1]$. (a) At $s=0$, the circuit consists of two layers. Between $s=0$ and $s=1/2$ the gates $u^{\dagger}_{x,x+1}(v_{0}^{\gamma},t)$ in the top layer (red) are swapped to the bottom layer while having their order of multiplication reversed. (b) The gate $u^{\dagger}_{0,1}(v_{0}^{\gamma},t)w_{0,1}(t)$ $u^{\dagger}_{-1,0}(v_{0}^{\gamma},t)w_{-1,0}(t)$] acts as the identity on site 0 for $t\in\{0,1\}$. It can thus be swapped (over $s\in[\tfrac{1}{2},\tfrac{3}{4}]$) so that it acts on site 11 [1-1] and its ancilla while ensuring that the whole circuit is still a loop. (c) Between $s=\tfrac{3}{4}$ and $s=1$ the circuit (red) is “rolled up”, such that it acts as the identity (gray dots) in any finite region.
@@
[…]
both act as the identity on site $x=0$ at $t=1$. Thus, in $s\in[\tfrac{1}{2},\tfrac{3}{4}]$, we can introduce $\mathrm{SWAP}(4(s-\tfrac{1}{2}))$ gates to the positions shown in Fig. 6(b) which “roll up” the circuit—the gates previously acting on site 0 now act on-site at $x=\pm 1$ between ancillae. Explicitly, for $s\in[\tfrac{1}{2},\tfrac{3}{4})$, the gate in $E(\gamma,t,s)$ acting across the $(00^{\prime},11^{\prime})$ edge is
@@
As $w^{\prime}_{0,1}(1)$ acts as the identity on site 0, the introduced swap gates cancel at $t=1$, and the entire circuit remains a loop for all $s$. By $s=\tfrac{3}{4}$, the gate acts as the identity on site 0 for all $t\in[0,1]$. We repeat an analogous deformation on the other gate with support on site 0. The part of the circuit which is supported on $x=0$ is thus rolled up to act on the ancillae at sites $x=\pm 1$.
This rolling up procedure can be iterated. At $s=\frac{3}{4}$, we see that the gate $w^{\prime}_{1,2}(t)w^{\prime}_{1^{\prime},1}(t)$ is the identity on sites $1$ and $1^{\prime}$ when $t=1$. Thus, by adding ancillae to site $2$, this gate can also be rolled up within $s\in[\tfrac{3}{4},\tfrac{7}{8}]$ [Fig. 6(c)]. We continue rolling up sites $x$ and x-x within $s\in[1-2^{-x-1},1-2^{-x-2}]$, such that, as $s\to 1$, we have $E(\gamma,t,s)\to\mathbbm{1}\otimes S(v_{0}^{\gamma},t)$ on any finite patch of the lattice. Technically, this makes $E(\gamma,t,s)$ continuous in the strong operator topology.